Project R delegates all the development work to outside companies. The estimated cash flows for project R are: (last column indicates expenditure) Beginning of year 1 150,000 dollars Contractor fees Beginning of year 2 250,000 dollars Contractor fees Beginning of year 3 250,000 dollars Contractor fees End of year 3 1,000,000 dollars Sales Table 1 Project S carries out all the development work in-house by purchasing the necessary equipment and using the company’s own staff. The cash flows for the project S are: (last column indicates expenditure) Beginning of year 1 325,000 dollars New equipment Throughout year 1 75,000 dollars Staff cost Throughout year 2 90,000 dollars Staff cost Throughout year 3 120,000 dollars Staff cost End of year 3 1,000,000 dollars Sales Table 2 The staff cost can be estimated to be paid uniformly throughout the year.

b) Find the IRR for the project R 


1
Expert's answer
2021-05-10T13:44:19-0400

"NPV=\sum_{\frac{FV}{(1+IRR)^1}+\frac{FV}{(1+IRR)^2}+\frac{FV}{(1+IRR)^3}}-IC=0"


"\sum_{\frac{-150 000}{(1+IRR)^1}+\frac{-250 000}{(1+IRR)^2}+\frac{750 000}{(1+IRR)^3}}-0=0"


IRR=0.2941 or 29.41%